What happens if you try to pay less than $5 per mo

How to do this question using while loop only?
Answer #1

The Fourier rework of the first signal is:

X(jω)=∫∞−∞x(t)e−jωtdt
X(jω)=∫−∞∞x(t)e−jωtdt
We take NN samples from x(t)x(t), and people samples will be denoted as x[0]x[0], x[1]x[1],...,x[n]x[n],...,x[N−1]x[N−1].

Since we tend to might assume every sample x[n]x[n] as Associate in Nursing impulse that has a district of x[n]x[n]:

X(jω)=∫(N−1)T0x(t)e−jωtdt
X(jω)=∫0(N−1)Tx(t)e−jωtdt
=x[0]e−j0+x[1]e−jωT+...+x[n]e−jωnT+...+x[N−1]e−jω(N−1)T
=x[0]e−j0+x[1]e−jωT+...+x[n]e−jωnT+...+x[N−1]e−jω(N−1)T
So, X(jω)X(jω) becomes

X(jω)=∑n=0N−1x[n]e−jωnT
X(jω)=∑n=0N−1x[n]e−jωnT
Since there ar solely a finite variety of computer file, the DFT treats the information as if it were amount, and evaluates the equation for the basic frequency:

ω=0,2πNT,2πNT×2,...,2πNT×n,...,2πNT×(N−1)
ω=0,2πNT,2πNT×2,...,2πNT×n,...,2πNT×(N−1)
Therefore, the separate Fourier rework of the sequence x[n]x[n] will be outlined as:

X[k]=∑n=0N−1x[n]e−j2πkn/N(k=0:N−1)
X[k]=∑n=0N−1x[n]e−j2πkn/N(k=0:N−1)
The equation will be written in matrix form:

⎡⎣⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢X[0]X[1]X[2]...X[N−1]⎤⎦⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥=⎡⎣⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢1111...11WW2W3...WN−11W2W4W6...W2(N−1)1W3W6W9...W3(N−1)..................1WN−1W2(N−1)W3(N−1)...W(N−1)(N−1)⎤⎦⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎡⎣⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢⎢x[0]x[1]x[2]...x[N−1]⎤⎦⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥⎥
[X[0]X[1]X[2]...X[N−1]]=[1111...11WW2W3...WN−11W2W4W6...W2(N−1)1W3W6W9...W3(N−1)..................1WN−1W2(N−1)W3(N−1)...W(N−1)(N−1)][x[0]x[1]x[2]...x[N−1]]
where W=e−j2π/NW=e−j2π/N and W=W2N=1W=W2N=1.


Quite a few individuals use WNWN for WW.

So, our final DFT equation will be outlined like this:

X[k]=∑n=0N−1x[n]WknN (k=0:N−1)
X[k]=∑n=0N−1x[n]WNkn (k=0:N−1)
The inverse of it:

x[n]=1N∑n=0N−1X[k]W−knN (k=0:N−1)